OpenStudy (anonymous):

lim (-3x+1)/sqrt(x^2+x) as x approaches -infinity

OpenStudy (anonymous):

eyeball it

OpenStudy (anonymous):

ignore everything but the leading terms \[\frac{-3x}{\sqrt{x^2}}=\frac{-3x}{|x|}\] as \(x \) goes to \(-\infty\) you get 3

OpenStudy (anonymous):

why do you ignore the other stuff?

OpenStudy (anonymous):

because is \(x\to \infty\) they are irrelevant

OpenStudy (anonymous):

think for example of \[\lim_{x\to \infty}\frac{2x^2+x+7}{x^2+5x}\] if say \[x=1000\] which is not a particularly large number, you have, in the numerator, a 2 in the millions place, some other stuff, in the denominator you have a 1 in the millions place, and some other stuff ratio is essentially 2 , the other stuff is in much lower places

OpenStudy (anonymous):

Okay, thank you!